# Sparse Feature Coactivation Reveals Causal Semantic Modules in Large Language Models

### v2

@@ Abstract
We identify semantically coherent, context-consistent network components in large language models (LLMs) using coactivation of sparse autoencoder (SAE) features collected from just a handful of prompts. Focusing on concept-relation prediction tasks, we show that ablating these components for concepts (e.g., countries and words) and relations (e.g., capital city and translation language) changes model outputs in predictable ways, while amplifying these components induces counterfactual responses. Notably, composing relation and concept components yields compound counterfactual outputs. Further analysis reveals that while most concept components emerge from the very first layer, more abstract relation components are concentrated in later layers. Lastly, we show that extracted components more comprehensively capture concepts and relations than individual features while maintaining specificity. Overall, our findings suggest a modular organization of knowledge accessed through compositional operations, and advance methods for efficient, targeted LLM manipulation.
@@ 1 Introduction
Sparse autoencoders (SAEs; Huben et al., 2024) recently emerged as a powerful tool for extracting interpretable, often monosemantic features from the internal representations of LLMs. SAEs achieve this by learning to reconstruct dense, entangled activations at each neural network layer through sparse codes, where individual features are incentivized to activate rarely and selectively. Despite this progress, it remains unclear how features from different layers organize and coordinate to produce responses (e.g., how an LLM may access and compose its encoded knowledge to respond “Beijing” to a prompt for China’s capital city). Cross-layer transcoders extend SAEs to approximate layerwise computations and globally interpret information flow in LLMs as computation graphs (Dunefsky et al., 2024; Ameisen et al., 2025), these graphs consist of hundreds of densely connected nodes, requiring manual feature grouping for interpretability. As such, there remains a need for a global, sparse, feature-level interpretation of LLMs’ relational knowledge and reasoning.
@@
In this work, we employ SAE feature coactivation to construct inter-layer feature networks for individual prompts about various factual and linguistic relational reasoning tasks. By filtering out denser features that activate across diverse contexts, we identify semantically coherent connected components within these networks representing specific concepts and relations across different prompts. We show that selective ablation and amplification of these components changes model outputs in predictable ways, overriding in-prompt instructions. In particular, composing a relation component with a concept component produces compound counterfactual outputs. As exemplified in Figure 1, a prompt about China’s capital activates components related to China and capital cities, eliciting the correct response of “Beijing” from the model. However, ablating the China component while amplifying the Nigeria component effectively steers its response to “Abuja”, whereas ablating the capital component while amplifying the language component steers it to “Chinese.” Furthermore, simultaneously ablating the China and capital components while amplifying the Nigeria and language components changes the model output to “English.”
In addition to contributing a sparse, modular structure that causally accounts for LLM relational reasoning processes, these components readily enable fine-grained interpretation. For example, they reveal the prominence and topology of specific concepts and relations; by ablating individual component features, we discover that country components tend to emerge from the very first network layer. In contrast, more abstract components for facts about countries are concentrated in later layers. Even within individual country fact components, we observe that nodes from deeper layers tend to have a stronger causal impact on model outputs. This aligns with our understanding of artificial neural networks and biological systems, where higher-level abstractions emerge in later processing stages. Lastly, our analysis demosntrates that our components offer a more comprehensive account of LLM relational reasoning than SAE features alone, and that components for each concept-relation pair are fairly specific to the functions we associate them with. Together, these results offer insights into the organization of relational knowledge in LLMs and establish sparse feature coactivation as an effective, lightweight approach for mechanistic interpretability and targeted model manipulation.
@@ 2 Methods
We run each input prompt through the model integrated with pre-trained SAEs [^2] We use SAE Lens to apply the width_16k/canonical variant from the gemma-scope-2b-pt-res-canonical release (Lieberum et al., 2024; Bloom et al., 2024). Each SAE maps the residual stream activation at layer $\ell$, $x_{\ell}\in\mathbb{R}^{d_{\text{model}}}$, to a sparse representation $\phi_{\ell}\in\mathbb{R}^{d_{\text{sae}}}$ where $d_{\text{sae}}=16384$. This produces an activation tensor $\Phi_{\ell}\in\mathbb{R}^{T\times d_{\text{sae}}}$ for each layer $\ell$, where $T$ is the number of non-BOS tokens in the prompt.
@@
We construct a directed graph $G=(V,E)$ where each node $(\ell,i)\in V$ corresponds to a selected feature $i\in S_{\ell}$. Edges $E$ connect nodes in adjacent layers according to the temporal correlation of their activation patterns across tokens in the prompt. Specifically, for feature pairs in adjacent layers $i\in S_{\ell}$ and $j\in S_{\ell+1}$, we compute the Pearson correlation coefficient $\rho$. A directed edge $e=((\ell,i),(\ell+1,j))$ is added to $E$ if $\rho(\Phi_{\ell}[:,i],\Phi_{\ell+1}[:,j])>\tau_{\text{corr}}=0.9$. Edge weights are assigned as $w(e)=\min(1.0,\rho)$.
@@
Some SAE features activate frequently across unrelated contexts, making them overly generic and hard to interpret. To reduce noise, we prune the graph using activation density scores from Neuronpedia (Lin, 2023). For each node $(\ell,i)\in V$, we retrieve its activation density $d_{\ell,i}$, defined as the fraction of tokens in a large corpus where the feature activates. We retain only sparse features, i.e., those with $d_{\ell,i}\leq\tau_{\text{density}}=0.01$, [^3] This follows Neuronpedia’s standard, which classifies features below this density as sparse and interpretable. creating a pruned graph $G_{\text{sparse}}$. We also remove any isolated nodes from $G_{\text{sparse}}$. We then use a straightforward BFS-based method implemented by NetworkX [^4] https://networkx.org/ to identify weakly connected components within $G_{\text{sparse}}$.
@@
To evaluate the functional significance of each component, we perform targeted interventions using TransformerLens (Nanda and Bloom, 2022). Specifically, we ablate or amplify the activations of SAE features in a given component during the model’s forward pass and measure the resulting shift in the probability distribution over next-token predictions. We quantify this shift using KL divergence between the original and perturbed distributions. A component is considered causal if its manipulation leads to systematic changes in model behavior, measured in Section 3.
Figure 2: We extract components from LLM queries to translate love into Spanish, French, and German. Each component is plotted with its feature count on the $x$-axis and the KL divergence between pre- and post-ablation output token distributions on the $y$-axis. Notably, for love and other concept, only a small number of components exert significant causal effects. For love, the top three components for each relation consistently correspond to either the queried word or language.
@@ 3 Experimental Results
For our initial experiments, we consider 4 multi-relational prediction tasks where the base model achieves 100% accuracy: (1) capital city, currency, and language of 10 countries, (2) translation of 11 words from English into Spanish, French, and German, and (3) synonym, antonym, past tense, capitalization, and first letter of 8 English verbs. [^5] Latter 2 tasks adapted in part from data used in Todd et al. (2024). See Appendix B for full specifications of tasks.
To collect model activations, we used simple prompt templates with in-context examples, e.g., for each country-capital pair, we used: “The capital city of Peru is Lima. The capital city of South Korea is Seoul. The capital city of Saudi Arabia is Riyadh. The capital city of {country} is”. Similar templates were used across tasks (Appendix C).
@@
Capital Spanish Capitalize
@@
Abuja 85. New 7.7 Lagos 17. rojo 38. La 4.1 Red 29. Like 21. Capital 5.6 like 12.
Lagos 11. Islamabad 7.3 the 12. Rojo 34. " 3.5 red 16. like 12. The 4.7 6.1
.38 Kathmandu 6.0 Nigeria 7.1 rojo 2.1 Bol 3.2 The 6.4 Capital 8.6 Yes 4.5 capital 5.6
Table 1: Top 3 output tokens and their likelihoods for various concept-relation pairs, before and after ablating concept or relation components. Underscore prefixes of tokens omitted for space. More examples in Appendix H.1.
@@ 3.1 Component Identification
To identify semantically coherent representative components for each concept and relation from these most impactful components, we applied a few approaches. First, we inspected their associated feature descriptions from Neuronpedia. [^7] These descriptions from Neuronpedia were generated by prompting GPT-4o mini (OpenAI et al., 2024) to summarize the texts that activate each feature over a large corpus. In many cases, the features within components had thematically coherent descriptions referring to common concepts or relations, e.g., most country fact task components (Appendix E).
However, these descriptions were often unreliable. For instance, none of the components obtained for prompts about Spain mention Spain in their feature descriptions, and some components for the word beautiful included features about Scotland. Additionally, some components for translation languages were described to be about programming. This suggests that even SAE features suffer from a degree of polysemanticity, possibly due to unavoidable limitations in the corpora and batching strategies used for training SAEs. Notably, this shows that it is impossible to consistently extract comparable components by simply performing a keyword search over SAE feature descriptions.
To address this in the country facts task, we performed component ablations and observed the resulting changes in the model’s top predicted tokens, selecting components that yielded the most compositional changes. As an example, the left 3 columns of Table 1 present results from ablation experiments using selected components obtained from prompts for the capital of China and Nigeria. Promisingly, when country components were ablated, the model’s top predicted tokens shifted predictably to the capitals of other countries. When fact components were ablated, the model assigned higher probabilities to country names.
For the remaining tasks, this approach was unsuccessful, possibly due to more complex composition operations at play than can be captured by a single component for a concept or relation. [^8] For example, steering with a particular component for here caused the model to output common locations for the prompted word, e.g., a word meaning ocean for the prompted word fish in both Spanish and French, a nontrivial operation. In these tasks, we instead represented each concept and relation by the union of multiple seemingly relevant causally impactful components. As shown in the other columns of Table 1, ablating these combined components for words/verbs in these tasks similarly caused the model to output other words in the prompted language of Spanish, or other capitalized words. Meanwhile, ablating relation components caused the model to repeat the input word.
@@ 3.2 Component Steering
Having identified distinct components for each concept and relation, we next investigate whether they can be used to steer model outputs individually and in combination. [^9] See Appendix G for details on steering strength selection. To test whether components generalize across different contexts, we applied a zero-shot prompt template different from the one used to collect the initial activations, e.g., for translation: “Q: What is the {Spanish/French/German} word for ’{word}’? Answer directly (two words max). A:”. We used comparable prompts for other tasks (Appendix C). Table 2 summarizes the results by task and relation, while Appendix I further breaks down performance by concepts.
Language, CN French, love Past Tense, break
$\cdot$, NG Capital, $\cdot$ Currency, $\cdot$ $\cdot$, red Spanish, $\cdot$ German, $\cdot$ $\cdot$, like Capitalize, $\cdot$ 1st Letter, $\cdot$
_English 71. _Beijing 96. _Yuan 15. _rouge 44. _amor 41. _Liebe 60. _liked 43. _Past 17. _B 24.
_Yoruba 5.6 _Be 1.5 _Ren 10. _Rouge 30. _el 9.5 _die 4.0 _Liked 43. _Has 7.8 _( 8.2
_Ha 4.8 _Peking .83 _RMB 6.8 _Sang 2.5 _Amor 8.5 _„ 3.3 _L 1.6 _Kick 4.4 _Past 5.8
@@
By ablating an in-prompt concept component and amplifying a target concept component, we successfully directed the LLM to respond to questions about relations for the in-prompt concept with counterfactual answers (i.e., from applying the relation to the target concept which is not actually queried in the prompt) 48-96% of the time (Table 2). For example, as shown in Table 3, the model consistently disregarded prompts’ references to China, love, and break, instead outputting counterfactual answers “English” (Nigeria’s language), “rouge” (red in French), and “liked” (like’s past tense). Even when incorrect answers were among top tokens, they were often relevant to the in-prompt relation and target concept, e.g., the model ranked “Yoruba” and “Ha[[usa]],” (other commonly spoken languages in Nigeria) just behind the correct answer of “English,” while it ranked a French word for blood just behind words for the target concept of red. This confirms that our identified concept components encode concept-specific information that causally determines model outputs.
We observed the most failures in the verb transformations task. Notably, some verbs had much lower concept steering success rate than others, e.g., possess and sink, which had respective concept steering success rates of about 6% and 20%. We observe that these verbs are also highly represented as in-prompt verbs in failures to steer to other target verbs. This suggests that our method did not yield stable components for these particular verbs, causing both failures to ablate and steer based on them. Interestingly, we see when steering to sink from break and hide, the model respectively generates sought and slid, possibly suggesting that components for sink may instead represent words that begin with s. Meanwhile, the capitalize relation had the lowest success rate; we observed that the model usually output the target word, but not capitalized. For hide, the model instead output the target verb in all capital letters. This may suggest that components for specific words can override prompt instructions when used for steering.
@@
By ablating an in-prompt relation while amplifying a target relation, we also successfully directed the model to respond to queries about the in-prompt relation as though they concerned the target relation 26-93% of the time (Table 2). As shown in Table 3, the model disregarded prompts’ references to country language, French translation, and past tense transformation, and produced correct counterfactual answers for other relations within each task, e.g., answering prompts to translate love into French with Spanish and German words for love. Even when incorrect answers were among top tokens, they were often relevant, e.g., other common languages or words in the target language. The only failure was the target transformation of capitalize, which yielded capitalized words other than the in-prompt verb break.
Unsurprisingly, the most failures again occurred in this task, specifically for the synonym, antonym, and past tense transformations. For the former two, this may be expected. Unlike the capital city or currency of a country, synonym and antonym do not have objective answers, and actually depend on a variety of finer-grained dimensions, e.g., word sense and part of speech. As observed in previous visualized token distributions after ablation and steering, the country fact and word translation components have clear roles in model outputs: to promote tokens related to specific countries or words, or that are relevant to a target fact (languages, currencies, or cities) or in a target language. However, a mechanism for synonym or antonym relations should not be expected to work like this, as many tokens could be a synonym or antonym, and reasonable answers thus depend more heavily on context. While past tense forms of verbs are more consistent, this relation is much broader and less specific than relations in other tasks, as every verb has a past tense form, and thus an LLM having a component to promote past tense inflections may be inefficient. Such relations are likely better captured in attention heads, which aligns with Todd et al. (2024)’s finding that antonym and past tense are modeled well by function vectors.
Currency, China German, love Antonym, break
Capital, Nigeria Language, Nigeria Spanish, red French, red Synonym, like Capitalize, like
_Abuja 70. _English 93. _rojo 45. _rouge 31. _love 46. _Opposite 9.3
_Lagos 26. _French 1.7 _el 9.8 _le 11. _loved 6.2 _Dis 7.9
_ 2.3 _Yoruba 1.5 _roja 4.2 _la 4.9 _lot 6.0 _ 6.3
@@
We conducted composite steering experiments where both concept and relation components were manipulated at once. By ablating both the in-prompt concept and in-prompt relation components while amplifying the target concept and target relation components, it was possible to steer the model to ignore both the in-prompt concept and in-prompt relation and answer about a different concept-relation pair 22-92% of the time (Table 2). Table 4 provides specific examples of composite steering success. As shown in the first 2 columns, when we ablated both the China and capital components while amplifying the Nigeria and currency components, the model answered “Naira” despite being asked about China’s capital. In the middle 2 columns, we see similar success in steering the model’s outputs to Spanish and French words for red, despite being prompted for the German translation of love. In the last 2 columns, when prompted for an antonym of break, we successfully steer the model to output a synonym of like, while steering it to output like capitalized (i.e., “Like”) was less successful, with the correct answer only ranked fourth among possible next tokens. Nonetheless, these findings demonstrate the significant composability of our components.

### v3

@@ Abstract
We identify semantically coherent, context-consistent network components in large language models (LLMs) using coactivation of sparse autoencoder (SAE) features collected from just a handful of prompts. Focusing on concept-relation prediction tasks, we show that ablating these components for concepts (e.g., countries and words) and relations (e.g., capital city and translation language) changes model outputs in predictable ways, while amplifying these components induces counterfactual responses. Notably, composing relation and concept components yields compound counterfactual outputs. Further analysis reveals that while most concept components emerge from the very first layer, more abstract relation components are concentrated in later layers. Lastly, we show that extracted components more comprehensively capture concepts and relations than individual features while maintaining specificity. Overall, our findings suggest a modular organization of knowledge and advance methods for efficient, targeted LLM manipulation. [^1] Code available at https://github.com/shanestorks/SAE-Semantic-Modules.
@@ 1 Introduction
Sparse autoencoders (SAEs; Huben et al., 2024) recently emerged as a powerful tool for extracting interpretable, often monosemantic features from the internal representations of LLMs. SAEs achieve this by learning to reconstruct dense, entangled activations at each neural network layer through sparse codes, where individual features are incentivized to activate rarely and selectively. Despite this progress, it remains unclear how features from different layers organize and coordinate to produce responses (e.g., how an LLM may access and compose its encoded knowledge to respond “Beijing” to a prompt for China’s capital city). Cross-layer transcoders (Dunefsky et al., 2024) extend SAEs by learning to approximate MLP computations and have been used to construct computation graphs that trace information flow across LLM layers Ameisen et al. (2025). However, these graphs can contain hundreds of densely connected nodes, making them difficult to interpret without manual grouping of features, motivating a more automated approach.
@@
In this work, we employ SAE feature coactivation to construct inter-layer feature networks for individual prompts about various factual and linguistic relational reasoning tasks. By filtering out features that activate across diverse contexts, we identify semantically coherent connected components representing specific concepts and relations. These components are consistent across prompts, and manipulating them predictably alters model outputs. As exemplified in Figure 1, a prompt about China’s capital activates components related to China and capital cities, eliciting the correct response of “Beijing” from the model. However, ablating features in the China component while amplifying those in the Nigeria component effectively steers its response to “Abuja”, whereas ablating the capital component while amplifying the language component steers it to “Chinese.” Furthermore, simultaneously ablating the China and capital components while amplifying the Nigeria and language components changes the model output to “English.” These components also readily visualize the topology of stored concepts and relations across layers. Our experiments yield the following findings:
  • Selective ablation and amplification of components consistently overrides prompted instructions, steering model outputs to predictable counterfactual answers in up to 98% of cases when steering based on individual concepts or relations, and up to 90% of cases when based on compound concept-relation pairs.
  • Conceptual components representing information about specific entities and words tend to emerge from the very first network layer, while more abstract relational components (e.g., for properties of countries or translation languages) concentrate in later layers. [^2] This aligns with our understanding of artificial neural networks and biological systems, where higher-level abstractions emerge in later processing stages DiCarlo et al. (2012); LeCun et al. (2015).
  • Our components offer a more comprehensive causal account of LLM relational reasoning than individual SAE features, and are mostly highly specific to the concepts and relations we associate them with (i.e., ablating them often preserves performance in other contexts).
Together, these results offer insights into the organization of relational knowledge in LLMs and establish sparse feature coactivation as an effective, lightweight approach for mechanistic interpretability and targeted model manipulation.
@@ 2 Methods
We run each input prompt through the LLM integrated with pre-trained SAEs. [^4] We use SAE Lens to apply the width_16k/canonical variant from the gemma-scope-2b-pt-res-canonical release (Lieberum et al., 2024; Bloom et al., 2024). Each SAE maps the residual stream activation at layer $\ell$, $x_{\ell}\in\mathbb{R}^{d_{\text{model}}}$, to a sparse representation $\phi_{\ell}\in\mathbb{R}^{d_{\text{sae}}}$ where $d_{\text{sae}}=16384$. This produces an activation tensor $\Phi_{\ell}\in\mathbb{R}^{T\times d_{\text{sae}}}$ for each layer $\ell$, where $T$ is the number of non-BOS tokens in the prompt.
@@
We construct a directed graph $G=(V,E)$ where each node $(\ell,i)\in V$ corresponds to a selected feature $i\in S_{\ell}$. Edges $E$ connect nodes in adjacent layers according to the temporal correlation of their activation patterns across tokens in the prompt. Specifically, for feature pairs in adjacent layers $i\in S_{\ell}$ and $j\in S_{\ell+1}$, we compute the Pearson correlation coefficient $\rho$. A directed edge $e=((\ell,i),(\ell+1,j))$ is added to $E$ if $\rho(\Phi_{\ell}[:,i],\Phi_{\ell+1}[:,j])>\tau_{\text{corr}}=0.9$.
@@
Some SAE features activate frequently across unrelated contexts, making them overly generic and hard to interpret. To reduce noise, we prune the graph using activation density scores from Neuronpedia (Lin, 2023). For each node $(\ell,i)\in V$, we retrieve its activation density $d_{\ell,i}$, i.e., the fraction of tokens in a large corpus where its corresponding feature activates. We retain only sparse features, i.e., those with $d_{\ell,i}\leq\tau_{\text{density}}=0.01$, [^5] This follows Neuronpedia’s standard, which classifies features below this density as sparse and interpretable. creating a pruned graph $G_{\text{sparse}}$. We also remove any isolated nodes from $G_{\text{sparse}}$. We then use a straightforward BFS-based method from NetworkX [^6] https://networkx.org/ to identify weakly connected components within $G_{\text{sparse}}$.
@@
To evaluate the functional significance of each component, we perform targeted interventions using TransformerLens (Nanda and Bloom, 2022). Specifically, we ablate (i.e., set to zero) the activations of SAE features in a given component during the LLM’s forward pass, [^7] When manipulating SAE features like this, we replace the LLM’s internal activations at each layer with the corresponding SAE decoder’s output after features are manipulated, then proceed with the forward pass., then measure the resulting shift in the probability distribution over next-token predictions. We quantify this shift using KL divergence between the original and perturbed distributions. A component is considered causal if its ablation causes a relatively high shift to other components and leads to systematic changes in model behavior, explored in Section 3.1 through various multi-relational prediction tasks.
We further validate components in Section 3.2, where we attempt to steer model behavior. Specifically, we ablate components associated with prompted information, then amplify one or more other components unrelated to the prompt, i.e., raise the activation for each SAE feature in the component by a proportion of its maximum activation observed during activation collection. We refer to this proportion as a steering strength $\alpha$. Steering is successful when it causes the model to generate a counterfactual response consistent with amplified component(s) rather than components associated with prompted information.
Figure 2: We extract components from LLM queries to translate love into Spanish, French, and German. Each component is plotted with its feature count on the $x$-axis and the total KL divergence between pre- and post-ablation output token distributions on the $y$-axis. For these and most other concept-relation pairs, only a few components exert significant causal effects. For love, the top three components for each language consistently correspond to either the queried word or language.
@@ 3 Experimental Results
We consider 3 multi-relational prediction tasks where the model is queried to generate a response (e.g., Beijing) that requires composing a given concept (e.g., China) and relation (e.g., capital city): [^8] Latter 2 tasks adapted in part from data used in Todd et al. (2024). See Appendix B for full specifications of tasks.
  1. 1.
    Country facts: capital city, currency, and language (relations) of 10 countries (concepts)
  2. 2.
    Word translation: translation of 11 English words (concepts) into Spanish, French, and German (relations)
  3. 3.
    Verb transformation: synonym, antonym, past tense, capitalization, and first letter (relations) of 8 English verbs (concepts)
Gemma 2 2B achieves 100% accuracy on each task. To collect model activations, we used simple prompt templates with in-context examples, e.g., for each country-capital pair, we used: “The capital city of Peru is Lima. The capital city of South Korea is Seoul. The capital city of Saudi Arabia is Riyadh. The capital city of {country} is”. Similar templates were used across tasks (Appendix C).
@@
Capital, $\frac{\text{China}}{\text{Nigeria}}$ Spanish, $\frac{\text{love}}{\text{red}}$ Capitalize, $\frac{\text{break}}{\text{like}}$
@@
Shanghai .23 Paris 6.1 Shanghai 5.0 amo 5.1 Malo 3.2 " 5.3 Capital 5.7 I 3.6 capital 4.3
Xi .11 Berlin 5.0 a 2.7 " 2.9 la 2.5 LOVE 2.7 The 5.0 Yes 2.3 The 3.3
Abuja 85. New 7.7 Lagos 17. rojo 38. La 3.5 Red 27. Like 21. Capital 5.6 like 12.
Lagos 11. Islamabad 7.3 the 12. Rojo 34. 3.4 red 15. like 12. The 4.7 6.1
.38 Kathmandu 6.0 Nigeria 7.1 rojo 2.1 " 3.2 " 8.1 Capital 8.6 Yes 4.5 capital 5.6
Nigeria .29 Delhi 5.7 called 5.1 Ro 2.0 El 2.8 The 5.1 LIKE 4.3 No 4.1 the 4.6
.27 Tehran 4.9 Abuja 3.4 Roja 1.9 Bol 2.4 4.1 The 4.0 3.3 Like 3.8
Table 1: Top 5 output tokens and their likelihoods for various concept-relation pairs, before and after ablating concept or relation components. Underscore prefixes of tokens omitted for space. More examples in Appendix H.1.
@@ 3.1 Component Identification
Given several impactful components, we then identify semantically coherent representative components for each concept and relation, i.e., components that yield predictable counterfactual outputs when manipulated (like in Figure 1). We explored a few approaches. First, we inspected their associated feature descriptions from Neuronpedia. [^10] These descriptions from Neuronpedia were generated by prompting GPT-4o mini (OpenAI et al., 2024) to summarize the texts that activate each feature over a large corpus. In many cases, the features within components had thematically coherent descriptions referring to common concepts or relations, e.g., most country fact task components (Appendix E).
However, these descriptions were often unreliable. For instance, none of the components obtained for prompts about Spain mentioned Spain in their feature descriptions, and some components for the word beautiful included features about Scotland. Additionally, some components for translation languages were described to be about programming. This suggests that even SAE features suffer from some polysemanticity. Notably, this shows that it is impossible to consistently extract comparable components by simple keyword search over SAE feature descriptions, necessitating methods like ours to group features based on task-specific activations.
To address this in the country facts task, we ablated components and observed the resulting changes in the model’s top predicted tokens, selecting those that yielded the most compositional changes. As an example, the left 3 columns of Table 1 present results of ablating selected components obtained from prompts for the capital of China and Nigeria. Promisingly, when ablating country components, the model’s top predicted tokens shifted to the capitals of other countries. When ablating fact components, the model assigned higher probabilities to country names.
For the remaining tasks, this approach was unsuccessful, possibly due to more complex composition operations at play than can be captured by a single component for a concept or relation. In these tasks, we instead represented each concept and relation by the union of multiple seemingly relevant causally impactful components. As shown in the other columns of Table 1, ablating these combined components for words/verbs in these tasks similarly caused the model to output other words in the prompted language of Spanish, or other capitalized words. Meanwhile, ablating relation components caused the model to repeat the input word.
@@ 3.2 Component Steering
Having identified distinct components for each concept and relation, we next investigate whether they can be used to steer model outputs individually and in combination. [^11] See Appendix G for details on steering strength tuning. To test whether components generalize across different contexts, we applied a zero-shot prompt template different from the one used to collect the initial activations, e.g., for translation: “Q: What is the {Spanish/French/German} word for ’{word}’? Answer directly (two words max). A:”. We used comparable prompts for other tasks, as listed in Appendix C.
Table 2 summarizes the results by task and relation, while Appendix I further breaks down performance by concepts, and Appendix J discusses common and interesting steering failure cases.
Language, CN French, love Past Tense, understand
$\cdot$, NG Capital, $\cdot$ Currency, $\cdot$ $\cdot$, red Spanish, $\cdot$ German, $\cdot$ $\cdot$, break Capitalize, $\cdot$ 1st Letter, $\cdot$
_English 71. _Beijing 96. _Yuan 12. _rouge 44. _amor 41. _Liebe 68. _broke 28. _Understand 6.7 _U 39.
_Yoruba 5.6 _Be 2.0 _Ren 8.4 _Rouge 30. _Amor 13. _die 2.4 _Broke 21. _Know 6.4 _understood 9.3
_Ha 4.8 _Peking .93 _China 6.1 _Sang 2.5 _el 6.9 _„ 2.1 _BRO 8.5 _Past 5.9 _Understand 8.5
_Nigeria 4.4 _BE .53 _Chinese 5.9 _Rou 1.8 _" 6.1 _lieben 1.7 _Bre 6.3 _Answer 5.7 _u 3.6
_Igbo 3.4 Beijing .37 _RMB 5.9 _ 1.2 _A 3.1 _Ich 1.6 _brake 2.6 _ 4.5 _United 3.2
@@
By ablating an in-prompt concept component and amplifying a target concept component, we successfully directed the LLM to respond to questions about relations for the in-prompt concept with counterfactual answers (i.e., from applying the relation to the target concept which is not actually queried in the prompt) 48-96% of the time on average (Table 2). For example, as shown in Table 3, the model consistently disregarded prompts’ references to China, love, and understand, instead outputting counterfactual answers “English” (Nigeria’s language), “rouge” (red in French), and “broke” (break’s past tense). Even when incorrect answers were among top tokens, they were often relevant to the in-prompt relation and target concept, e.g., the model ranked “Yoruba,” “Ha[[usa]],” and “Igbo” (other commonly spoken languages in Nigeria) just behind the correct answer of “English,” while it ranked a French word for blood just behind words for the target concept of red. This confirms that our identified concept components encode concept-specific information that causally determines model outputs.
We observed the most failures in the verb transformations task. Notably, some verbs had a much lower concept steering success rate than others, e.g., possess (6%) and sink (20%). These same verbs caused failures in both directions: steering to them (as target verbs) and steering away from them (as in-prompt verbs). This suggests that our method did not yield stable components for these particular verbs. Interestingly, we saw that when steering to sink from break and hide under the past tense relation, the model respectively generated “sought” and “slid”, possibly suggesting that components for sink may instead represent words that begin with s. Meanwhile, the capitalize relation had the lowest success rate; we observed that the model usually output the target word, but not capitalized. For hide, the model instead output the target verb in all capital letters. This may suggest that components for specific verbs can override prompt instructions when used for steering, perhaps indicating a lower degree of composability than other tasks.
Currency, China German, love Antonym, break
Capital, Nigeria Language, Nigeria Spanish, red French, red Synonym, like 1st Letter, like
_Abuja 70. _English 96. _rojo 53. _rouge 45. _love 56. _like 31.
_Lagos 25. _Yoruba 1.1 _el 7.2 _le 6.7 _loved 8.5 _Like 20.
_ 3.3 _French .74 _roja 5.6 _Rouge 4.2 _few 3.8 _L 6.3
_Nigeria .34 _Spanish .60 _" 3.1 _ 3.7 _loves 3.6 _ant 3.2
_… .32 _Igbo .54 _El 2.2 _la 3.4 _lot 3.4 _Ant 2.5
@@
By ablating an in-prompt relation while amplifying a target relation, we also successfully directed the model to respond to queries about the in-prompt relation as though they concerned the target relation 23-98% of the time (Table 2). As shown in Table 3, the model disregarded prompts’ references to country language, French translation, and past tense transformation, and produced correct counterfactual answers for other relations within each task, e.g., answering prompts to translate love into French with Spanish and German words for love. Even when incorrect answers were among top tokens, they were often relevant, e.g., words in the target language, or “Know” for the capitalized form of understand (a synonym for the correct answer). In some cases, correct answers competed with answers related to the prompt context, suggesting that some in-prompt information remains activated. For example, “Chinese” appeared among the top tokens when steering to the currency relation even after the in-prompt component for the language relation was ablated. Similarly, “understood” appeared among top tokens when steering to first letter after the in-prompt component for past tense was ablated.
Unsurprisingly, the most failures again occurred in this task, specifically for the synonym, antonym, and past tense transformations. For the former two, this may be expected. Unlike the capital city or currency of a country, synonym and antonym do not have objective answers, and actually depend on a variety of finer-grained dimensions, e.g., word sense and part of speech. As observed in previous visualized token distributions after ablation and steering, the country fact and word translation components have clear roles in model outputs: to promote tokens related to specific countries or words, or that are relevant to a target fact (languages, currencies, or cities) or in a target language. However, a mechanism for synonym or antonym relations should not be expected to work like this, as many tokens could be a synonym or antonym, and reasonable answers thus depend more heavily on context. While past tense forms of verbs are more consistent, this relation is much broader and less specific than relations in other tasks, as every verb has a past tense form, and thus an LLM learning a specialized module to promote past tense inflections may be inefficient. Such relations are thus likely better captured in attention heads, which aligns with Todd et al. (2024)’s finding that antonym and past tense are modeled well by function vectors.
@@
We conducted composite steering experiments where both concept and relation components were manipulated at once. By ablating both the in-prompt concept and in-prompt relation components while amplifying the target concept and target relation components, it was possible to steer the model to ignore both the in-prompt concept and in-prompt relation and answer about a different concept-relation pair 19-90% of the time (Table 2). Table 4 provides specific examples of composite steering success. As shown in the first 2 columns, when we ablated both the China and capital components while amplifying the Nigeria and currency components, the model answered “Naira” despite being asked about China’s capital. In the middle 2 columns, we see similar success in steering the model’s outputs to Spanish and French words for red, despite being prompted for the German translation of love. In the last 2 columns, when prompted for an antonym of break, we successfully steer the model to output a synonym of like, while steering it to output the first letter of like was less successful, with the correct answer only ranked third after “like” and “Like”. We also observed interesting failure cases where in-prompt information intermingled with target information, e.g., steering from the antonym of break to include capitalized yielded “Exclude,” a capitalized antonym of include. Nonetheless, these findings demonstrate significant composability of many components.